OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (4, –6); m = 3/5

OpenStudy (anonymous):

need help to find this answer

OpenStudy (anonymous):

I point slope form is y=mx+b and what you've got is (x,y) m=slope then all you have to do is put it in order.

OpenStudy (anonymous):

ohh ok so would the answer be Y- 4= 3/5 (x+6) ????

OpenStudy (anonymous):

Good try but it would be more like -6=3/5+4

OpenStudy (anonymous):

Y-6 = 3/5 (X + 4) OR Y +6 = 3/5 (X-4)

OpenStudy (anonymous):

Since you don't need to change any signs the first one would be correct. You don't need the y at the beginning because the (-6) is a y coordinates.

OpenStudy (anonymous):

ok that makes sense now :) thank you

OpenStudy (anonymous):

No problem